Show that there is a 96.7°angle between the vectors 5a +37 and 3ā – 5 if ā= -i+j – k and 7 = 3i + 2j – 3k .
Answer #1

Solution:- Let the vectoes are vi = 5a + 3 b V2 = 34 b a - ā=-i+jok & b = 31 toj 1 3K Н VI = 5(-itj - k) + 3( 31 +23-3k ) = -il -13 (:1333 = 37x9 = 3√37 537 3837 ) -13 + cos 2 13 c d = car-l (-13 T d ñ 96.72575224° ㅋ e 96.7° T Angle between two vecto